Dimensions of a Rectangle A rectangle is 5 feet
longer than it is wide and has an area of 126 square
feet. What are its dimensions?

Answer by oberobic(2304)

Area = Length * Width
A = L * W
.
L = W+5
.
L*W = 126
.
(W+5)*W = 126
.
W^2 +5W - 126 = 0
.
(W+14)(W-9) = 0
.
W = -14 or 9.
.
A negative width is irrelevant, so we go with W = 9.
.
L = 9+5
L = 14
.
Done.